Find the area included between the curve x^3 + y^3 = 3axy and its asymptote
1
Expert's answer
2020-04-15T17:33:35-0400

"x^3+y^3=3axy"

"[1+(\frac{y}{x})^3]=\frac{3a}{x}(\frac{y}{x})"

"\frac{[1+(\frac{y}{x})^3]}{(\frac{y}{x})}=\frac{3a}{x}"

As x tends to "\infty" and "-\infty" ; RHS tends to 0,so LHS also tend to 0.

So , "1+(\frac{y}{x})^3" tends to 0,thus,"\frac{y}{x}=-1"

Hence the asymptote has slope −1 and so is of the form "y+x=k"

Substituting "y=k-x" in the equation of curve.

"\u22123k\u00b2x+3kx\u00b2\u2212k\u00b3 = 3ax(k-x)"

"\u22123k\u00b2\/x+3k\u2212k\u00b3\/x\u00b2 = 3a[(k\/x)-1 ]"

"k=-a"

This,equation of asymptote is "x+y+a=0"

The product of slopes of the line symmetrical to "y=x" and the equation of asymptote "x+y+a=0" is (-1) i.e. they are at right angle to each other.Whence if the axes of references are turned through an angle of 45° ,the new X axis coincides with the symmetrical line "y=x" with a changed equation of curve having asymptote parallel to the new Y-axis instead the oblique asymptote "x+y+a=0" .


After transformation new X and Y are:

"X=xcos(45°)-ysin(45°)=\frac{x-y}{\sqrt{2}}"

"Y=ysin(45°)+ycos(45°)=\frac{x+y}{\sqrt{2}}"

"x=\frac{X+Y}{\sqrt{2}};y=\frac{Y-X}{\sqrt{2}}"

New equation of curve becomes

"(\frac{X+Y}{\sqrt{2}})^3+(\frac{Y-X}{\sqrt{2}})^3=3a(\frac{X+Y}{\sqrt{2}})(\frac{Y-X}{\sqrt{2}})"

"\frac{1}{2\sqrt{}2}((X+Y)^3+(Y-X)^3)=(\frac{3a}{2})(Y^2-X^2)"

"(X^3+3XY^2)=\frac{3a}{\sqrt{2}}(Y^2-X^2)"

"3Y^2(X+\frac{a}{\sqrt{2}})-X^2(\frac{3a}{\sqrt{2}}-X)=0"

"3Y^2(X+b)-X^2(3b-X)=0" where "b=\frac{a}{\sqrt{2}}"

Clearly,the asymptote of the equation parallel to Y-axis is X=-b.

Hence,the area between the curve and its asymptote is given by

"A=2\int_{-b}^{0}YdX"

"A=2\int_{-b}^{0}\frac{-x}{\sqrt{3}}(\frac{3b-x}{x+b})^{\frac{1}{2}}dX"

"A=\frac{-2}{\sqrt{3}}\int_{-b}^{0}(\frac{x(3b-x)}{\sqrt{(x+b)(3b-x)}})dX"

Putting "X=b(1-2cos\theta);dx=2bsin\theta d \theta" for "X=-b,\theta=0;X=0,\theta=\frac{π}{3}"

"A=\frac{-2}{\sqrt{3}}\int_0^{\frac{π}{3}}\frac{b(1-2cos\theta)(3b-(-b-2bcos\theta)).2bsin\theta}{\sqrt{((b-2bcos\theta)+b)(3b-(b-2bcos\theta))}}d\theta"

"A=\frac{-2}{\sqrt{3}}\int_0^{\frac{π}{3}}2b^2(1-cos\theta-2cos^2\theta)d\theta"

"A=\frac{4b^2}{\sqrt{3}}\int_0^{\frac{π}{3}}(cos\theta+cos2\theta)d\theta"

"A=\frac{4b^2}{\sqrt{3}}[(sin\theta+\frac{sin2\theta}{2})]_0^\frac{π}{3}"

"A=3b^2=\frac{3}{2}a^2"